# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 4. Main results
The first main result asserts the local strong well-posedness of the interaction problem for initial data in $\mathrm{X}_{\gamma}$. In contrast to [19] in the weak setting, we consider here the general $Q$-tensor model with arbitrary ratio $\xi$ of tumbling and alignment effects in the strong setting and especially obtain uniqueness.
@@
$h^{\prime}$ $\in\mathrm{W}^{1,p}(0,T^{\prime})^{3},\enspace h\in\mathrm{L}^{\infty}(0,T^{\prime})^{3},\enspace\Omega\in\mathrm{W}^{1,p}(0,T^{\prime})^{3}\enspace\text{and}\enspace\pi^{\mathcal{F}}\in\mathrm{L}^{p}\bigl{(}0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}(\cdot))\cap\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot))\bigr{)}.$
@@ 5. Transformation to a fixed domain
(5.1) $\partial_{t}X_{0}(t,y)&=m(t)(X_{0}(t,y)-h(t))+h^{\prime}(t),&&\enspace\text{on}\enspace(0,T)\times\mathbb{R}^{3},\\ X_{0}(0,y)&=y,&&\enspace\text{for}\enspace y\in\mathbb{R}^{3}.\end{aligned}\right.$
@@
(5.2) $\partial_{t}X(t,y)&=b(t,X(t,y)),&&\enspace\text{on}\enspace(0,T)\times\mathbb{R}^{3},\\ X(0,y)&=y,&&\enspace\text{for}\enspace y\in\mathbb{R}^{3}.\end{aligned}\right.$
@@
Besides, the Picard-Lindelöf theorem yields the existence of a unique solution $X\in\mathrm{C}^{1}\bigl{(}(0,T);\mathrm{C}^{\infty}(\mathbb{R}^{3})^{3}\bigr{)}$ to (5.2) for given $h^{\prime}$, $\Omega\in\mathrm{W}^{1,p}(0,T)^{3}$. This solution has continuous mixed partial derivatives. From the uniqueness of the solution, we deduce the bijectivity of the function $X(t,\cdot)$, and its inverse is denoted by $Y(t,\cdot)$. Thanks to $\operatorname{div}b=0$, Liouville’s theorem yields that $X$ and $Y$ are both volume-preserving, $\mathrm{J}_{X}(t,y)\mathrm{J}_{Y}(t,X(t,y))=\operatorname{Id}_{3}$ and $\det\mathrm{J}_{X}(t,y)=\det\mathrm{J}_{Y}(t,x)=1$, with $\mathrm{J}_{X}$ and $\mathrm{J}_{Y}$ representing the Jacobian matrices of $X$ and $Y$ as usual. The inverse $Y$ of $X$ solves $\partial_{t}Y(t,x)=b^{(Y)}(t,Y(t,x))$ on $(0,T)\times\mathbb{R}^{3}$, where the right-hand side is given by $b^{(Y)}(t,y)\coloneqq-\mathrm{J}_{X}^{-1}(t,y)b(t,X(t,y))$, and $Y(0,x)=x$. Let us stress that $b^{(Y)}$ and $Y$ possess the same space and time regularity as $b$ and $X$. Moreover, the interval of existence of the solution is restricted by the condition $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in(0,T)$.
@@
$v(t,y)$ $\coloneqq\mathrm{J}_{Y}(t,X(t,y))u^{\mathcal{F}}(t,X(t,y)),\enspace Q(t,y)\coloneqq Q^{\mathcal{F}}(t,X(t,y)),\enspace\ell(t)\coloneqq\mathbb{O}(t)^{\top}h^{\prime}(t),\enspace\omega(t)\coloneqq\mathbb{O}(t)^{\top}\Omega(t),$
$p(t,y)$ $\coloneqq\pi^{\mathcal{F}}(t,X(t,y))\enspace\text{and}\enspace\Sigma(v(t,y),p(t,y),Q(t,y))\coloneqq\mathbb{O}(t)^{\top}\tilde{\Sigma}(\mathbb{O}(t)v(t,y),p(t,y),Q(t,y))\mathbb{O}(t).$
@@
(5.3) $g^{ij}=\sum_{k=1}^{3}(\partial_{k}Y_{i})(\partial_{k}Y_{j}),\enspace g_{ij}=\sum_{k=1}^{3}(\partial_{i}X_{k})(\partial_{j}X_{k})\enspace\text{and}\enspace\Gamma^{i}_{jk}=\frac{1}{2}\sum_{l=1}^{3}g^{il}(\partial_{k}g_{jl}+\partial_{j}g_{kl}-\partial_{l}g_{jk}).$
From [18, (3.13)] and [8, (3.13)], we recall the transformed Laplacians
(5.4) $(\mathcal{L}_{1}v)_{i}$ $=\sum_{j,k=1}^{3}\partial_{j}(g^{jk}\partial_{k}v_{i})+2\sum_{j,k,l=1}^{3}g^{kl}\Gamma_{jk}^{i}(\partial_{l}v_{j})+\sum_{j,k,l=1}^{3}\Bigl{(}\partial_{k}(g^{kl}\Gamma^{i}_{kl})+\sum_{m=1}^{3}g^{kl}\Gamma^{m}_{jl}\Gamma^{i}_{km}\Bigr{)}v_{j}\enspace\text{and}\enspace$
@@
$(\mathcal{M}v)_{i}=\sum\limits_{j=1}^{3}(\partial_{t}Y_{j})(\partial_{j}v_{i})+\sum\limits_{j,k=1}^{3}\bigl{(}\Gamma^{i}_{jk}(\partial_{t}Y_{k})+(\partial_{k}Y_{i})(\partial_{j}\partial_{t}X_{k})\bigr{)}v_{j},\enspace(\mathcal{N}(v))_{i}=\sum\limits_{j=1}^{3}v_{j}(\partial_{j}v_{i})+\sum\limits_{j,k=1}^{3}\Gamma^{i}_{jk}v_{j}v_{k}$
and $(\mathcal{G}p)_{i}=\sum\limits_{j=1}^{3}g^{ij}(\partial_{j}p)$.
@@
(5.5) $[B_{\sigma}(Q)A]_{i}\coloneqq\sum_{j,k=1}^{3}\Bigl{(}\bigl{(}(\partial_{j}Q_{ik})(\Delta A_{kj})+Q_{ik}(\partial_{j}\Delta A_{kj})\bigr{)}-\bigl{(}(\partial_{j}\Delta A_{ik})Q_{kj}+(\Delta A_{ik})(\partial_{j}Q_{kj})\bigr{)}\Bigr{)}.$
The next term to be addressed is $\tau_{\mathrm{h}}$ from (3.1). We further split it into $\tau_{\mathrm{h}}\coloneqq\xi(\tau_{\mathrm{h},1}+\tau_{\mathrm{h},2}+\tau_{\mathrm{h},3})$, where $\tau_{\mathrm{h},1}\coloneqq-\nicefrac{{2}}{{3}}H_{\mathrm{h}}$, $\tau_{\mathrm{h},2}\coloneqq-\{Q,H_{\mathrm{h}}\}$ and $\tau_{\mathrm{h},3}\coloneqq 2(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(QH_{\mathrm{h}})$. We also introduce $B_{\tau_{\mathrm{h},i}}(Q)A$ such that $B_{\tau_{\mathrm{h},i}}(Q)Q=\operatorname{div}\tau_{\mathrm{h},i}$ for $i=1,2,3$. The first term is $[B_{\tau_{\mathrm{h},1}}A]_{i}\coloneqq-\nicefrac{{2}}{{3}}\sum_{j=1}^{3}\partial_{j}(\Delta A_{ij}-A_{ij})$. Observing that $\{Q,H_{\mathrm{h}}\}=Q(\Delta Q)+(\Delta Q)Q-2Q^{2}$, we define
(5.6) $[B_{\tau_{\mathrm{h},2}}(Q)A]_{i}\coloneqq-(B_{\sigma,1}(Q)A)_{i}+(B_{\sigma,2}(Q)A)_{i}+2\sum_{j,k=1}^{3}\bigl{(}(\partial_{j}A_{ik})Q_{kj}+Q_{ik}(\partial_{j}A_{kj})\bigr{)}.$
@@
To capture $-\nabla Q\odot\nabla Q$, we set $[B_{\tau_{\mathrm{l},2}}(Q)A]_{i}\coloneqq-\sum_{j,k=1}^{2}\bigl{(}(\partial_{i}Q_{jk})\Delta A_{kj}+\sum_{l=1}^{3}(\partial_{l}Q_{kj})(\partial_{l}\partial_{i}A_{jk})\bigr{)}$. The term $B_{\tau_{\mathrm{l},3}}(Q)A$ associated to $\tau_{\mathrm{l},3}$ is of a similar form as $B_{\tau_{\mathrm{l},1}}(Q)A$.
@@
(5.7) $\mathcal{B}_{\sigma}(Q)A\coloneqq\mathcal{B}_{\sigma,1}(Q)A+\mathcal{B}_{\sigma,2}(Q)A\coloneqq\mathcal{B}_{\sigma,11}(Q)A+\mathcal{B}_{\sigma,12}(Q)A+\mathcal{B}_{\sigma,2}(Q)A,\enspace\text{where}\enspace$
$(\mathcal{B}_{\sigma,11}(Q)A)_{i}$ $\coloneqq\sum_{j,k=1}^{3}(\mathcal{L}_{2}A)_{kj}\sum_{l=1}^{3}(\partial_{l}Q_{ik})(\partial_{j}Y_{l}),\enspace(\mathcal{B}_{\sigma,12}(Q)A)_{i}\coloneqq\sum_{j,k=1}^{3}Q_{ik}\Biggl{(}\sum_{l=1}^{3}\biggl{(}\sum_{m=1}^{3}\Bigl{(}\sum_{n=1}^{3}(\partial_{n}\partial_{m}\partial_{l}A_{kj})$
@@
$\mathcal{B}_{\tau_{\mathrm{h}}}(Q)A$ $\coloneqq\xi(\mathcal{B}_{\tau_{\mathrm{h},1}}A+\mathcal{B}_{\tau_{\mathrm{h},2}}(Q)A+\mathcal{B}_{\tau_{\mathrm{h},3}}(Q)A),\enspace\mathcal{B}_{\tau_{\mathrm{l}}}(Q)A\coloneqq\xi\mathcal{B}_{\tau_{\mathrm{l},1}}(Q)A+\mathcal{B}_{\tau_{\mathrm{l},2}}(Q)A+\xi\mathcal{B}_{\tau_{\mathrm{l},3}}(Q)A,$
$\mathcal{B}_{\tau}(Q)A$ $\coloneqq\mathcal{B}_{\tau_{\mathrm{h}}}(Q)A+\mathcal{B}_{\tau_{\mathrm{l}}}(Q)A\enspace\text{and}\enspace\mathcal{B}_{\tau}(Q)\coloneqq\mathcal{B}_{\tau}(Q)Q\enspace\text{as well as}\enspace\mathcal{B}_{\sigma}(Q)\coloneqq\mathcal{B}_{\sigma}(Q)Q.$
@@
The term $H(Q)$ from (2.3) transforms to $\mathcal{H}(Q)=\mathcal{L}_{2}Q-Q+(Q^{2}-\operatorname{tr}(Q^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}})-\operatorname{tr}(Q^{2})Q$. Then the complete transformed system associated to (2.5)–(2.9) on the fixed domain $(0,T)\times\mathcal{F}_{0}$ takes the shape
(5.11) $\partial_{t}v+(\mathcal{M}-\mathcal{L}_{1})v+\mathcal{N}(v)+\mathcal{G}p&=\mathcal{B}_{\tau}(Q)+\mathcal{B}_{\sigma}(Q),\enspace\operatorname{div}v=0,&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}Q}+(\partial_{t}Y\cdot\nabla)Q+(v\cdot\nabla)Q-\mathcal{S}(v,Q)&=\mathcal{H}(Q),&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}\ell^{\prime}+m_{\mathcal{S}}(\omega\times\ell)&=-\int\limits_{\partial\mathcal{S}_{0}}\Sigma(v,p,Q)n\,\mathrm{d}\Gamma,&&\enspace\text{in}\enspace(0,T),\\ J_{0}\omega^{\prime}-\omega\times(J_{0}\omega)&=-\int\limits_{\partial\mathcal{S}_{0}}y\times\Sigma(v,p,Q)n\,\mathrm{d}\Gamma,&&\enspace\text{in}\enspace(0,T).\end{aligned}\right.$
Here $n=\mathbb{O}^{\top}(t)\nu(t)$ is the unit outward normal to $\partial\mathcal{F}_{0}$. The boundary and initial conditions are
(5.12) $v=0,\enspace\partial_{\nu}Q=0\enspace\text{on}\enspace(0,T)\times\partial\mathcal{O},\enspace v=\ell+\omega\times y,\enspace\partial_{\nu}Q=0\enspace\text{on}\enspace(0,T)\times\partial\mathcal{S}_{0}\enspace\text{and}\enspace$
(5.13) $v(0)=v_{0},\enspace Q(0)=Q_{0},\enspace\ell(0)=\ell_{0}\enspace\text{and}\enspace\omega(0)=\omega_{0}.$
@@
$v$ $\in\mathrm{W}^{1,p}\bigl{(}0,T^{\prime};\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\bigr{)}\cap\mathrm{L}^{p}\bigl{(}0,T^{\prime};\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\bigr{)},\enspace Q\in\mathrm{W}^{1,p}\bigl{(}0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr{)}\cap\mathrm{L}^{p}\bigl{(}0,T^{\prime};\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr{)},$
$\ell$ $\in\mathrm{W}^{1,p}(0,T^{\prime})^{3},\enspace\omega\in\mathrm{W}^{1,p}(0,T^{\prime})^{3}\enspace\text{and}\enspace p\in\mathrm{L}^{p}\bigl{(}0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})\bigr{)}.$
@@
Let $p>4$, let $\mathcal{O}\subset\mathbb{R}^{3}$ be a bounded domain of class $\mathrm{C}^{3}$, and consider the domains of the rigid body and the fluid at time zero $\mathcal{S}_{0}$ and $\mathcal{F}_{0}$. Besides, assume that $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$ for some $r>0$. Then there exists $\eta_{0}>0$ such that for all $\eta\in(0,\eta_{0})$, there are $\delta_{0}>0$, $C>0$ so that for every $\delta\in(0,\delta_{0})$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\overline{B}_{\mathrm{X}_{\gamma}}(0,\delta)$, it holds that $X\in\mathrm{C}^{1}([0,\infty);\mathrm{C}^{2}(\mathbb{R}^{3})^{3})$, with $X(\tau,\cdot)\colon\mathcal{F}(\tau)\to\mathcal{F}_{0}$ being $\mathrm{C}^{2}$-diffeomorphisms for all $\tau\in[0,\infty)$, and (5.11)–(5.13) has a unique solution $(v,Q,\ell,\omega,p)$ with
@@
$+\|\mathrm{e}^{\eta(\cdot)}\ell\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\omega\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}p\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0}))}\leq C\delta.$

### v2

@@ 4. Main results
The first main result asserts the local strong well-posedness of the interaction problem for initial data in $\mathrm{X}_{\gamma}$. In contrast to [20], which establishes the existence of a weak solution to a slightly simplified version of the $Q$-tensor model, we consider here the general $Q$-tensor model with arbitrary ratio $\xi$ of tumbling and alignment effects in the strong setting, and we obtain uniqueness of the strong solution.
@@
$h^{\prime}$ $\in\mathrm{W}^{1,p}(0,T^{\prime})^{3},\kern 5.0pth\in\mathrm{L}^{\infty}(0,T^{\prime})^{3},\kern 5.0pt\Omega\in\mathrm{W}^{1,p}(0,T^{\prime})^{3}\kern 5.0pt\text{and}\kern 5.0pt\pi^{\mathcal{F}}\in\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}(\cdot))\cap\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot))\bigr).$
@@ 5. Transformation to a fixed domain
(5.1) $\left\{\begin{aligned} \partial_{t}X_{0}(t,y)&=m(t)(X_{0}(t,y)-h(t))+h^{\prime}(t),&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\mathbb{R}^{3},\\ X_{0}(0,y)&=y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathbb{R}^{3}.\end{aligned}\right$.
@@
(5.2) $\left\{\begin{aligned} \partial_{t}X(t,y)&=b(t,X(t,y)),&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\mathbb{R}^{3},\\ X(0,y)&=y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathbb{R}^{3}.\end{aligned}\right$.
@@
Besides, the Picard-Lindelöf theorem yields the existence of a unique solution $X\in\mathrm{C}^{1}\bigl((0,T);\mathrm{C}^{\infty}(\mathbb{R}^{3})^{3}\bigr)$ to (5.2) for given $h^{\prime}$, $\Omega\in\mathrm{W}^{1,p}(0,T)^{3}$. This solution has continuous mixed partial derivatives. From the uniqueness of the solution, we deduce the bijectivity of the function $X(t,\cdot)$, and its inverse is denoted by $Y(t,\cdot)$. Thanks to $\operatorname{div}b=0$, Liouville’s theorem yields that $X$ and $Y$ are both volume-preserving, $\mathrm{J}_{X}(t,y)\mathrm{J}_{Y}(t,X(t,y))=\operatorname{Id}_{3}$ and $\det\mathrm{J}_{X}(t,y)=\det\mathrm{J}_{Y}(t,x)=1$, with $\mathrm{J}_{X}$ and $\mathrm{J}_{Y}$ representing the Jacobian matrices of $X$ and $Y$ as usual. The inverse $Y$ of $X$ solves $\partial_{t}Y(t,x)=b^{(Y)}(t,Y(t,x))$ on $(0,T)\times\mathbb{R}^{3}$, where the right-hand side is given by $b^{(Y)}(t,y)\coloneqq-\mathrm{J}_{X}^{-1}(t,y)b(t,X(t,y))$, and $Y(0,x)=x$. Let us stress that $b^{(Y)}$ and $Y$ possess the same space and time regularity as $b$ and $X$. Moreover, the interval of existence of the solution is restricted by the condition $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in(0,T)$. In fact, we can write $\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq\|h(t)\|_{\mathbb{R}^{3}}+\|\mathbb{O}(t)-\mathbb{I}\|_{\mathbb{R}^{3\times 3}}|y|.$ As the rigid body moves with a continuous velocity, thanks to the estimates in Section 7, it is possible to find a sufficiently small time $T>0$ such that $\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq\frac{r}{2}$ is satisfied for large initial data. For small initial data, we will see that the exponential stability of the semigroup associated with the linearized problem at zero as revealed in Section 6 ensures that $\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq\frac{r}{2}$. Combining with the assumption $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$, we obtain $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in(0,\infty)$, so no collision occurs in finite time.
@@
$v(t,y)$ $\coloneqq\mathrm{J}_{Y}(t,X(t,y))u^{\mathcal{F}}(t,X(t,y)),\kern 5.0ptQ(t,y)\coloneqq Q^{\mathcal{F}}(t,X(t,y)),\kern 5.0pt\ell(t)\coloneqq\mathbb{O}(t)^{\top}h^{\prime}(t),\kern 5.0pt\omega(t)\coloneqq\mathbb{O}(t)^{\top}\Omega(t),$
$\pi(t,y)$ $\coloneqq\pi^{\mathcal{F}}(t,X(t,y))\kern 5.0pt\text{and}\kern 5.0pt\Sigma(v(t,y),\pi(t,y),Q(t,y))\coloneqq\mathbb{O}(t)^{\top}\tilde{\Sigma}(\mathbb{O}(t)v(t,y),\pi(t,y),Q(t,y))\mathbb{O}(t).$
@@
(5.3) $g^{ij}=\sum_{k=1}^{3}(\partial_{k}Y_{i})(\partial_{k}Y_{j}),\kern 5.0ptg_{ij}=\sum_{k=1}^{3}(\partial_{i}X_{k})(\partial_{j}X_{k})\kern 5.0pt\text{and}\kern 5.0pt\Gamma^{i}_{jk}=\frac{1}{2}\sum_{l=1}^{3}g^{il}(\partial_{k}g_{jl}+\partial_{j}g_{kl}-\partial_{l}g_{jk}).$
From [19, (3.13)] and [8, (3.13)], we recall the transformed Laplacians
(5.4) $(\mathcal{L}_{1}v)_{i}$ $=\sum_{j,k=1}^{3}\partial_{j}(g^{jk}\partial_{k}v_{i})+2\sum_{j,k,l=1}^{3}g^{kl}\Gamma_{jk}^{i}(\partial_{l}v_{j})+\sum_{j,k,l=1}^{3}\Bigl(\partial_{k}(g^{kl}\Gamma^{i}_{kl})+\sum_{m=1}^{3}g^{kl}\Gamma^{m}_{jl}\Gamma^{i}_{km}\Bigr)v_{j}\kern 5.0pt\text{and}\kern 5.0pt$
@@
$(\mathcal{M}v)_{i}=\sum\limits_{j=1}^{3}(\partial_{t}Y_{j})(\partial_{j}v_{i})+\sum\limits_{j,k=1}^{3}\bigl(\Gamma^{i}_{jk}(\partial_{t}Y_{k})+(\partial_{k}Y_{i})(\partial_{j}\partial_{t}X_{k})\bigr)v_{j},\kern 5.0pt(\mathcal{N}(v))_{i}=\sum\limits_{j=1}^{3}v_{j}(\partial_{j}v_{i})+\sum\limits_{j,k=1}^{3}\Gamma^{i}_{jk}v_{j}v_{k}$
and $(\mathcal{G}\pi)_{i}=\sum\limits_{j=1}^{3}g^{ij}(\partial_{j}\pi)$.
@@
(5.5) $B_{\sigma}(Q)A]_{i}\coloneqq\sum_{j,k=1}^{3}\Bigl(\bigl((\partial_{j}Q_{ik})(\Delta A_{kj})+Q_{ik}(\partial_{j}\Delta A_{kj})\bigr)-\bigl((\partial_{j}\Delta A_{ik})Q_{kj}+(\Delta A_{ik})(\partial_{j}Q_{kj})\bigr)\Bigr).$
The next term to be addressed is $\tau_{\mathrm{h}}$ from (3.1). We further split it into $\tau_{\mathrm{h}}\coloneqq\xi(\tau_{\mathrm{h},1}+\tau_{\mathrm{h},2}+\tau_{\mathrm{h},3})$, where $\tau_{\mathrm{h},1}\coloneqq-\nicefrac{{2}}{{3}}H_{\mathrm{h}}$, $\tau_{\mathrm{h},2}\coloneqq-\{Q,H_{\mathrm{h}}\}$ and $\tau_{\mathrm{h},3}\coloneqq 2(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(QH_{\mathrm{h}})$. We also introduce $B_{\tau_{\mathrm{h},i}}(Q)A$ such that $B_{\tau_{\mathrm{h},i}}(Q)Q=\operatorname{div}\tau_{\mathrm{h},i}$ for $i=1,2,3$. The first term is $B_{\tau_{\mathrm{h},1}}A]_{i}\coloneqq-\nicefrac{{2}}{{3}}\sum_{j=1}^{3}\partial_{j}(\Delta A_{ij}-A_{ij})$. Observing that $\{Q,H_{\mathrm{h}}\}=Q(\Delta Q)+(\Delta Q)Q-2Q^{2}$, we define
(5.6) $B_{\tau_{\mathrm{h},2}}(Q)A]_{i}\coloneqq-(B_{\sigma,1}(Q)A)_{i}+(B_{\sigma,2}(Q)A)_{i}+2\sum_{j,k=1}^{3}\bigl((\partial_{j}A_{ik})Q_{kj}+Q_{ik}(\partial_{j}A_{kj})\bigr).$
@@
To capture $\nabla Q\odot\nabla Q$, we set $B_{\tau_{\mathrm{l},2}}(Q)A]_{i}\coloneqq-\sum_{j,k=1}^{2}\bigl((\partial_{i}Q_{jk})\Delta A_{kj}+\sum_{l=1}^{3}(\partial_{l}Q_{kj})(\partial_{l}\partial_{i}A_{jk})\bigr)$. The term $B_{\tau_{\mathrm{l},3}}(Q)A$ associated to $\tau_{\mathrm{l},3}$ is of a similar form as $B_{\tau_{\mathrm{l},1}}(Q)A$.
@@
(5.7) $\mathcal{B}_{\sigma}(Q)A\coloneqq\mathcal{B}_{\sigma,1}(Q)A+\mathcal{B}_{\sigma,2}(Q)A\coloneqq\mathcal{B}_{\sigma,11}(Q)A+\mathcal{B}_{\sigma,12}(Q)A+\mathcal{B}_{\sigma,2}(Q)A,\kern 5.0pt\text{where}\kern 5.0pt$
$(\mathcal{B}_{\sigma,11}(Q)A)_{i}$ $\coloneqq\sum_{j,k=1}^{3}(\mathcal{L}_{2}A)_{kj}\sum_{l=1}^{3}(\partial_{l}Q_{ik})(\partial_{j}Y_{l}),\kern 5.0pt(\mathcal{B}_{\sigma,12}(Q)A)_{i}\coloneqq\sum_{j,k=1}^{3}Q_{ik}\Biggl(\sum_{l=1}^{3}\biggl(\sum_{m=1}^{3}\Bigl(\sum_{n=1}^{3}(\partial_{n}\partial_{m}\partial_{l}A_{kj})$
@@
$\mathcal{B}_{\tau_{\mathrm{h}}}(Q)A$ $\coloneqq\xi(\mathcal{B}_{\tau_{\mathrm{h},1}}A+\mathcal{B}_{\tau_{\mathrm{h},2}}(Q)A+\mathcal{B}_{\tau_{\mathrm{h},3}}(Q)A),\kern 5.0pt\mathcal{B}_{\tau_{\mathrm{l}}}(Q)A\coloneqq\xi\mathcal{B}_{\tau_{\mathrm{l},1}}(Q)A+\mathcal{B}_{\tau_{\mathrm{l},2}}(Q)A+\xi\mathcal{B}_{\tau_{\mathrm{l},3}}(Q)A,$
$\mathcal{B}_{\tau}(Q)A$ $\coloneqq\mathcal{B}_{\tau_{\mathrm{h}}}(Q)A+\mathcal{B}_{\tau_{\mathrm{l}}}(Q)A\kern 5.0pt\text{and}\kern 5.0pt\mathcal{B}_{\tau}(Q)\coloneqq\mathcal{B}_{\tau}(Q)Q\kern 5.0pt\text{as well as}\kern 5.0pt\mathcal{B}_{\sigma}(Q)\coloneqq\mathcal{B}_{\sigma}(Q)Q.$
@@
The term $H(Q)$ from (2.3) transforms to $\mathcal{H}(Q)=\mathcal{L}_{2}Q-Q+(Q^{2}-\operatorname{tr}(Q^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}})-\operatorname{tr}(Q^{2})Q$. Then for the unit outward normal $n=\mathbb{O}^{\top}(t)\nu(t)$ to $\partial\mathcal{F}_{0}$, which is thus directed inwards to $\partial\mathcal{S}_{0}$, the complete transformed system associated to (2.5)–(2.9) on the fixed domain $(0,T)\times\mathcal{F}_{0}$ takes the shape
(5.11) $\left\{\begin{aligned} \partial_{t}v+(\mathcal{M}-\mathcal{L}_{1})v+\mathcal{N}(v)+\mathcal{G}\pi&=\mathcal{B}_{\tau}(Q)+\mathcal{B}_{\sigma}(Q),\kern 5.0pt\operatorname{div}v=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}Q}+(\partial_{t}Y\cdot\nabla)Q+(v\cdot\nabla)Q-\mathcal{S}(v,Q)&=\mathcal{H}(Q),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}\ell^{\prime}+m_{\mathcal{S}}(\omega\times\ell)&=-\int\limits_{\partial\mathcal{S}_{0}}\Sigma(v,\pi,Q)n\,\mathrm{d}\Gamma,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}\omega^{\prime}-\omega\times(J_{0}\omega)&=-\int\limits_{\partial\mathcal{S}_{0}}y\times\Sigma(v,\pi,Q)n\,\mathrm{d}\Gamma,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T).\end{aligned}\right$.
The boundary and initial conditions are
(5.12) $v=0,\kern 5.0pt\partial_{n}Q=0\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0ptv=\ell+\omega\times y,\kern 5.0pt\partial_{n}Q=0\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0}\kern 5.0pt\text{and}\kern 5.0pt$
(5.13) $v(0)=v_{0},\kern 5.0ptQ(0)=Q_{0},\kern 5.0pt\ell(0)=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\omega(0)=\omega_{0}.$
@@
$v$ $\in\mathrm{W}^{1,p}\bigl(0,T^{\prime};\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\bigr)\cap\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\bigr),\kern 5.0ptQ\in\mathrm{W}^{1,p}\bigl(0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr)\cap\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr),$
$\ell$ $\in\mathrm{W}^{1,p}(0,T^{\prime})^{3},\kern 5.0pt\omega\in\mathrm{W}^{1,p}(0,T^{\prime})^{3}\kern 5.0pt\text{and}\kern 5.0pt\pi\in\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})\bigr).$
@@
Let $p>4$, let $\mathcal{O}\subset\mathbb{R}^{3}$ be a bounded domain of class $\mathrm{C}^{3}$, and consider the domains of the rigid body and the fluid at time zero $\mathcal{S}_{0}$ and $\mathcal{F}_{0}$. Besides, assume that $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$ for some $r>0$. Then there exists $\eta_{0}>0$ such that for all $\eta\in(0,\eta_{0})$, there are $\delta_{0}>0$, $C>0$ so that for every $\delta\in(0,\delta_{0})$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\overline{B}_{\mathrm{X}_{\gamma}}(0,\delta)$, it holds that $X\in\mathrm{C}^{1}([0,\infty);\mathrm{C}^{2}(\mathbb{R}^{3})^{3})$, with $X(\tau,\cdot)\colon\mathcal{F}(\tau)\to\mathcal{F}_{0}$ being $\mathrm{C}^{2}$-diffeomorphisms for all $\tau\in[0,\infty)$, and (5.11)–(5.13) has a unique solution $(v,Q,\ell,\omega,\pi)$ with
@@
$+\|\mathrm{e}^{\eta(\cdot)}\ell\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\omega\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\pi\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0}))}\leq C\delta.$